# Memory-Centric Embodied Question Answer

### v1

@@ Abstract
Embodied Question Answering (EQA) requires agents to autonomously explore and understand the environment to answer context-dependent questions. Existing frameworks typically center around the planner, which guides the stopping module, memory module, and answering module for reasoning. In this paper, we propose a memory-centric EQA framework named MemoryEQA. Unlike planner-centric EQA models where the memory module cannot fully interact with other modules, MemoryEQA flexible feeds memory information into all modules, thereby enhancing efficiency and accuracy in handling complex tasks, such as those involving multiple targets across different regions. Specifically, we establish a multi-modal hierarchical memory mechanism, which is divided into global memory that stores language-enhanced scene maps, and local memory that retains historical observations and state information. When performing EQA tasks, the multi-modal large language model is leveraged to convert memory information into the required input formats for injection into different modules. To evaluate EQA models’ memory capabilities, we constructed the MT-HM3D dataset based on HM3D, comprising 1,587 question-answer pairs involving multiple targets across various regions, which requires agents to maintain memory of exploration-acquired target information. Experimental results on HM-EQA, MT-HM3D, and OpenEQA demonstrate the effectiveness of our framework, where a 19.8% performance gain on MT-HM3D compared to baseline model further underscores memory capability’s pivotal role in resolving complex tasks.
Figure 1: This task requires the agent to identify multiple relevant entities involved in the question. Planner-centric method fail to recall that the living room sofa has already been found, and thus it searches for the sofa again, resulting in redundant exploration. In contrast, reasoning framework guided by memory chose an unexplored direction.
@@ 1 Introduction
Embodied Question Answering (EQA) [11, 14, 34, 16] is a challenging task in robotics, requiring agents to resolve questions via interaction within a dynamic and evolving environment. Existing methods [27, 39, 12, 7, 10] typically center around a planner (as shown in the left of Figure 2), coordinating modular components including memory module, answering module, and stopping module. The planner relies only on current observations to invoke the other modules. However, these methods often suffer from inefficient exploration and inaccurate response when faced with complex tasks (e.g., the tasks requires exploring multiple regions to gather information about numerous targets, while retaining memory of all targets referenced in the questions). The underlying reason is that performing complex tasks necessitates the retention of previously explored information across all modules, yet the planner-centric framework fails to facilitate direct interaction between the memory module and other modules. As shown in Figure 1, in the question “Are the sofas in the living room and the media room the same color?”, the agent must explore to multiple objects (“sofa in the living room”, “sofa in the media room”) for comparative reasoning. In the planner-centric framework, due to the lack of interaction between the memory module and planner, the effective information gathered during exploration remain invisible to planner. Specifically, after obtaining the knowledge that “the sofa in the living room is white”, the planner still inefficiently re-explore the same white sofa or fail the task entirely, as this critical input is unavailable during subsequent decision-making.
In this paper, we proposed MemoryEQA, a memory-centric EQA framework (as shown in the right of Figure 2) that maintains a memory database and infuses memory information into all modules, as shown in Figure 1. Unlike the planner-centric methods, where the planner relies only on semantic map, but semantic map is unable to feed into other modules, the proposed memory-centric framework injects memory information into various modules to guide reasoning, enabling efficient exploration and accurate responses.
To achieve this goal, we need to overcome two challenges when designing the memory module. (1) A single form of memory cannot meet the needs of different modules. For instance, the planner requires continuous semantic values (the use of high-dimensional embeddings to model semantic similarity through geometric proximity in a latent space) to guide the next planning objective, while the answering module necessitates observations and historical nature language dialogues pertinent to the query. (2) Natural language memory is difficult to integrate with semantic map and challenging to maintain. Semantic map is structured data (e.g., coordinates, attribute tables), whereas natural language exists as linear sequences, requiring complex conversions (such as mapping “sofa coordinates (x,y)” to “the white sofa on the right”).
To solve the above challenges, we design a multi-modal hierarchical memory, with a corresponding memory storage, retrieval, updating, and injection mechanism. Specifically, multi-modal hierarchical memory comprises global and local memory. Global memory is constructed by combining semantic map with information from object detectors, depth maps, and the agent’s current state, capturing entity locations and the agent’s observation. The local memory stores the current observations, captions of observations obtained through the vision-language model (VLM), and the agent’s current state. For the planner, we project the candidate directions for the next step into pixel space and annotate these candidate directions on the current RGB observation. The memory information is then incorporated as contextual prompts for the VLM to select an optimal direction from the candidates. Regarding the answering and stopping modules, we retrieve the most memory entries relevant to the given question from the memory bank and utilize them as contextual inputs for the VLM. Consequently, natural language serves as an intermediary to adapt memory information to different functional modules. Furthermore, during the construction of natural language-based memory, we employ a detector to localize objects within the scene, followed by leveraging the VLM to extract fine-grained descriptions of these objects. This approach not only enhances the semantic map with natural language annotations but also augments the natural language memory with world coordinate information, thereby enabling integration between natural language memory and the semantic map. Additionally, we have meticulously designed a memory update mechanism to ensure the real-time and relevance of the memory information.
@@
To evaluate memory ability of EQA models, we constructed the MT-HM3D dataset based on HM3D [26], including 1587 QA pairs. Specifically, we utilize the Habitat simulator [30] to sample 5 to 10 images per scene with constrained distances between sampling points. We then prompt GPT-4o [23] to generate questions about multiple targets across different regions, which require the agent to memorize all question-relevant information during exploration. Finally, low-quality data filtering is performed through GPT-4o-mini [23] and manual verification.
Our contributions are as follows:
@@
  • We propose a memory-centric EQA framework that provides a multi-modal hierarchical memory with a corresponding memory retrieval and injection mechanism, and a memory real-time update mechanism for efficient exploration and reasoning.
@@
  • We constructed the MT-HM3D dataset based on the HM3D, offering more complex QA pairs about multiple targets across different regions, effectively evaluating EQA models’ memory capabilities.
@@
  • MemoryEQA achieves an 18.9% improvement in success rate over the baseline model on the MT-HM3D dataset, demonstrating the importance of memory capability in solving complex tasks. Additionally, it sets a new state-of-the-art (SOTA) on the open-vocabulary EQA dataset OpenEQA and achieves comparable success rates to previous best methods on HM-EQA.
Figure 2: Comparison between the planner-centric framework and memory-centric framework.
@@ 2.1 Embodied Question Answering
Embodied question answering [11, 14, 39, 4, 12] has become a challenging task for testing the robot’s ability to autonomously plan tasks and establish semantic understanding of the environment. Das et al. [11] first proposed the EQA task and utilized a planner-controller navigation module to accomplish EQA task in virtual environments. Allen et al. [27] constructed a 2D semantic map of the environment for specific tasks to guide exploration. Answer et al. [2] and Xie et al. [37] developed offline embodied modules that could be queried by large language models (LLMs). Arjun et al. [21] utilized video memory to answer implicit questions using long-context VLMs. Saumya et al. [29] bridged the gap between semantic memory and planners by placing the planner within a compact scene representation. In our work, we position memory as the core module of the embodied question answering framework, enabling the planner, stopping criteria, and question-answering module to reason based on memory information. The memory-centric framework overcomes the limitations of planner-centric methods by infusing memory information into all modules, enabling more efficient exploration and accurate responses through enhanced inter-module interaction.
@@ 2.2 Memory Mechanisms in Embodied Agent
Early works utilized the hidden state of the LSTM [15] navigation model as a dense memory representation of the scene [30, 35]. However, the expressive power of this single state vector is limited, making it challenging to handle the storage of memory information in complex 3D real-world scenarios. Recent approaches have modeled scene memory as 2D metric grid [1, 3, 5, 6], topological map [28, 36], and 3D semantic map[9, 24, 31, 29] to serve as global memory information. Although these methods have shown promising performance in navigation task, the complexity of indoor environments, coupled with the lack of local memory information, makes it difficult for models to tackle fine-grained EQA effectively.
@@
Abrar et al. [2] used retrieval-augmented LLM-agent to retrieve relevant memories by forming function calls and answer questions based on a real-time memory-building process. Chen et al. [8] and Xie et al. [37] employed retrieval-augmented generation (RAG) to retrieve relevant images from accumulated observations and used a VLM for inference to generate responses. However, planning that does not rely on memory information will lead to repetitive exploration, resulting in the agent’s inability to answering questions correctly. To this end, we propose a memory-centric EQA framework, MemoryEQA, which maintains a hierarchical memory and flexibly injects memory information into various modules to achieve efficient exploration and accurate responses.
3 Memory-Centric EQA Framework
The overall framework is illustrated in Figure 3, with the memory module in its center, injecting memory information to other modules for reasoning. We define the agent observation sequence as $O=\{{o_{i}}\}_{i=0}^{N_{t}}$, state sequence as $S=\{{s_{i}}\}_{i=0}^{N_{t}}$, and decision sequence as $D=\{{d_{i}}\}_{i=0}^{N_{t}}$, where $N_{t}$ represents the total number of exploration steps. The input question is denoted as $q$. The hierarchical memory module is =$\mathcal{M}=\{M_{j}\}_{j=0}^{N_{s}}$, where $N_{s}$ is the number of scenes. The memory retriever is $\mathcal{F}_{r}$, the answering module is $\mathcal{F}_{a}$, the planner is $\mathcal{F}_{p}$, and the stopping module is $\mathcal{F}_{s}$.
@@ 3 Memory-Centric EQA Framework
Figure 3: The overall framework of memory-centric EQA. Given the input question and agent observation, we first update the hierarchical memory. Subsequently, question-relevant memories are retrieved through the memory retriever and injected into the respective modules. We concatenate the memories and the question, which are then fed into the planner to determine the next navigation point. Once the agent arrives at a navigation point, it assesses whether to halt its journey. If the decision is ‘No’, the process loops back to the memory update step; if ‘Yes’, the agent responds to the question.
We take the current observations and the question as input, first to update the memory, and then use $\mathcal{F}_{r}$ to retrieve the required memory information. Next, we utilize $\mathcal{F}_{s}$ to determine whether the exploration should be stopped. If sufficient information is available to answer the question, the retrieved memory, the question, and the current state are fed into $\mathcal{F}_{a}$ to generate a response. If the information is insufficient, the retrieved memory, question, state, and observations are input into the planner $\mathcal{F}_{p}$ to derive the next state $s_{i+1}$. In summary, we formulate our framework as
{$r&=\mathcal{F}_{a}(q,s_{t},\mathcal{F}_{r}(q,\mathcal{M})),\quad if\ \ \mathcal{F}_{s}(q,\mathcal{F}_{r}(q,\mathcal{M}))=\text{Yes},\\ s_{i+1}&=\mathcal{F}_{p}(s_{i},o_{i},q,\mathcal{F}_{r}(q,\mathcal{M})),\quad else,\end{aligned}\right.$ (1)
where $r$ represents the agent response and $s_{t}$ is the status when the agent stops exploring.
Figure 4: Hierarchical Memory. The global memory (yellow box) is constructed by annotating entity states on a real-time generated map, and local memory consist of he visual observation and object information (gray box). All memories are persisted and used as a database for scene retrieval (green box).
3.1 Planner
We employ a dynamic 3D voxel-based mapping approach for robotic exploration. The environment is represented using an extensible voxel grid with initial dimensions $L\times W\times H$, where height $H$ remains fixed at $3.5m$ (typical floor height) while length $L$ and width $W$ dynamically expand as exploration progresses. Each cubic voxel with edge length $l$ is updated in real-time through Truncated Signed Distance Function (TSDF) fusion[22, 40]: upon acquiring depth image, the system first updates occupancy states for all visible voxels while selectively marking exploration status only for voxels within a constrained field of view.
For path planning purposes, the 3D voxel map is projected onto a 2D point map M, where each point’s status is determined by two criteria: (1) it is marked as traversable when all voxels below camera elevation height are free space, and (2) considered explored only when all voxels along the vertical dimension have been explored. Building upon this representation, the system implements frontier-based exploration [38] for autonomous navigation by detecting boundaries between explored and unexplored regions, sampling candidate poses from these frontiers, and determining optimal exploration directions to achieve efficient environment mapping.
3.2 Hierarchical Memory
As shown in Figure 4, our hierarchical memory is divided into local memory and global memory. The corresponding memory storage, retriever, injector, and memory update mechanism collectively form a closed loop of “construction-utilization-updating”.
Local Memory.
Local memory is composed of visual observations, goal descriptions, planning information, and the agent’s current state. Specifically, based on the current visual observation, we use an object detector to identify the objects present in the current view, and obtain fine-grained descriptions of these objects by prompting a VLM. Planning information includes the current planning step and the decision (derived from the offset between the navigation point and the current position). The agent’s current state encompasses its location, orientation, and a coarse-grained description of the space it is in (obtained by prompting the VLM for a description of the current observation).
Gloabl Memory.
Building upon the real-time constructed map, the global memory annotates room information and target information. Among these, room information primarily includes the room category and its location (represented by the position from which the agent observes the room). Target information encompasses the target’s position, category, description, as well as the position and orientation from which the agent observes the target. Details on prompts about memory construction can be found in the supplemental materials.
3.2.1 Memory Storage

### v2

@@ Abstract
Embodied Question Answering (EQA) requires agents to autonomously explore and comprehend the environment to answer context-dependent questions. Typically, an EQA framework consists of four components: a planner, a memory module, a stopping module, and an answering module. However, the memory module is utilized inefficiently in existing methods, as the information it stores is leveraged solely for the answering module. Such a design may result in redundant or inadequate exploration, leading to a suboptimal success rate. To solve this problem, we propose MemoryEQA, an EQA framework centered on memory, which establishes mechanisms for memory storage, update, and retrieval, allowing memory information to contribute throughout the entire exploration process. Specifically, we convert the observation into structured textual representations, which are stored in a vector library following a fixed structure. At each exploration step, we utilize a viewpoint comparison strategy to determine whether the memory requires updating. Before executing each module, we employ an entropy-based adaptive retrieval strategy to obtain the minimal yet sufficient memory information that satisfies the requirements of different modules. The retrieved module-specific information is then integrated with the current observation as input to the corresponding module. To evaluate EQA models’ memory capabilities, we constructed the benchmark based on HM3D called MT-HM3D, comprising 1,587 question-answer pairs involving multiple targets across various regions, which requires agents to maintain memory of exploration-acquired target information. Experimental results on HM-EQA, MT-HM3D, and OpenEQA demonstrate the effectiveness of our framework, where a 9.9% performance gain on MT-HM3D compared to baseline models further underscores the memory capability’s pivotal role in solving complex tasks.
@@ 1 Introduction
Embodied Question Answering (EQA) [das2018embodied, gordon2018iqa, wijmans2019embodied, jiang2025beyond] is a challenging task in robotics that requires agents to actively interact with dynamic and evolving environments to answer natural language questions. Existing EQA approaches [ren2024explore, yu2019multi, das2018neural, chen2023not, dai2024think, jiang2025beyond] are typically designed around a planner, which coordinates other modules including the memory, answering, and stopping modules. However, in these planner-centric methods, the memory module is often underutilized, whose information is only leveraged by the answering module, leading to inefficient exploration and inaccurate responses, thereby reducing overall task success rates. For example, given complex tasks that require exploring multiple regions to gather information for reasoning, an agent must effectively retain and share past observations across all modules to enable efficient planning and exploration. Yet, planner-centric designs typically rely only on the agent’s current observation for planning and stopping decisions, resulting in redundant or insufficient exploration. As illustrated in Figure LABEL:fig:case, when answering a question “Are the sofas in the living room and the media room the same color?”, the agent must explore two objects (“sofa in the living room,” “sofa in the media room”) to perform comparative reasoning. In the planner-centric framework, since the planner lacks direct access to the memory module, crucial information collected during exploration remains inaccessible. Consequently, even after discovering that “the sofa in the living room is white,” the planner may inefficiently re-explore the same sofa or fail to complete the task, as this critical knowledge is unavailable during subsequent decision-making.
In this paper, we propose MemoryEQA, a memory-centric EQA framework that maintains a memory library and infuses memory information into all modules to guide reasoning, enabling efficient exploration and accurate responses. Specifically, MemoryEQA consists of four components: a planner, a memory module, a stopping module, and an answering module. For the planner, we augment the widely used frontier-based planner with memory guidance, leveraging memory information to steer planning. The memory module is equipped with mechanisms for storage, update, and retrieval, allowing memory to guide all processes in EQA. For the stopping and answering modules, we use different natural language queries to retrieve relevant information from the memory module and perform reasoning with a multi-modal large language model.
In doing so, we need to address a key challenge in the memory module design, that is, the instability of long-term memory updates and retrievals in dynamic, complex environments. Irrelevant memories may be retrieved and adversely affect exploration, planning, reasoning, and answering. In other words, long-term memory in embodied environments is easily corrupted by noisy updates and irrelevant retrievals. To address this, we propose a viewpoint-contrastive memory update rule for updating natural language memories, and an entropy-based adaptive retrieval strategy to select the most relevant memory information to the user query, mitigating the impact of redundant or irrelevant information on subsequent reasoning.
@@
To evaluate the memory capability of EQA models, we built a benchmark based on HM3D [ramakrishnan2021habitat] with 1587 QA pairs, called MT-HM3D. Specifically, we utilize the Habitat simulator [szot2021habitat] to sample 5 to 10 images per scene with constrained distances between sampling points. We then prompt GPT-4o [gpt4o] to generate questions about multiple targets across different regions, which require the agent to memorize all question-relevant information during exploration. Finally, low-quality data filtering is performed through manual verification. Our contributions are as follows:
@@
  • We propose a memory-centric EQA framework that provides a multi-modal hierarchical memory with a viewpoint-contrastive memory update rule and an entropy-based adaptive retrieval strategy for efficient exploration and reasoning.
@@
  • We construct the benchmark MT-HM3D based on the HM3D, offering more complex QA pairs about multiple targets across different regions, effectively evaluating EQA models’ memory capabilities.
@@
  • MemoryEQA achieves a 9.9% improvement in success rate over baseline methods on MT-HM3D, demonstrating the importance of memory capability in solving complex tasks. Additionally, it sets a new state-of-the-art on the open-vocabulary EQA dataset OpenEQA and achieves comparable success rates to previous best methods on HM-EQA.
@@ 2.1 Embodied Question Answering
Embodied question answering [das2018embodied, gordon2018iqa, yu2019multi, cangea2019videonavqa, das2018neural] has become a challenging task for testing the robot’s ability to autonomously plan tasks and establish semantic understanding of the environment. Das et al. [das2018embodied] first proposed the EQA task and utilized a planner-controller navigation module to accomplish EQA task in virtual environments. Allen et al. [ren2024explore] constructed a 2D semantic map of the environment for specific tasks to guide exploration. Answer et al. [anwar2024remembr] and Xie et al. [xie2024embodied] developed offline embodied modules that could be queried by large language models (LLMs). Arjun et al. [majumdar2024openeqa] utilized video memory to answer implicit questions using long-context Visal Language Models (VLMs). Saumya et al. [saxena2024grapheqa] bridged the gap between semantic memory and planners by placing the planner within a compact scene representation. Different from existing methods, we position memory as the core module of the embodied question answering framework, enabling the planner, stopping criteria, and question-answering module to reason based on memory information. The memory-centric framework overcomes the limitations of planner-centric methods by infusing memory information into all modules, enabling more efficient exploration and accurate responses through enhanced inter-module interaction.
@@ 2.2 Memory Mechanisms in Embodied Agent
Early works utilized the hidden state of the LSTM [graves2012long] navigation model as a dense memory representation of the scene [szot2021habitat, wijmans2019dd]. However, the expressive power of this single state vector is limited, making it challenging to handle the storage of memory information in complex 3D real-world scenarios. Recent approaches have modeled scene memory as 2D metric grid [anderson2019chasing, blukis2018mapping, cartillier2021semantic, chaplot2020object], topological map [savinov2018semi, wu2019bayesian], and 3D semantic map[cheng2018geometry, prabhudesai2019embodied, tung2019learning, saxena2024grapheqa] to serve as global memory information. Although these methods have shown promising performance in navigation task, the complexity of indoor environments, coupled with the lack of local memory information, makes it difficult for models to tackle fine-grained EQA effectively.
@@
Abrar et al. [anwar2024remembr] used retrieval-augmented LLM-agent to retrieve relevant memories by forming function calls and answer questions based on a real-time memory-building process. Chen et al. [cheng2024efficienteqa] and Xie et al. [xie2024embodied] employed retrieval-augmented generation (RAG) to retrieve relevant images from accumulated observations and used a VLM for inference to generate responses. However, planning that does not rely on memory information will lead to repetitive exploration, resulting in the agent’s inability to answering questions correctly. To this end, MemoryEQA establishes a comprehensive memory mechanism encompassing storage, update, and retrieval, enabling memory information to more effectively facilitate reasoning for efficient exploration and accurate answering.
@@
Figure 1: The overall framework of memory-centric EQA. Given the input question and agent observation in multiple steps, we gradually update the structural memory. Subsequently, question-relevant memories are retrieved through the memory retriever and input into the respective modules. We concatenate the memories and the question, which are then fed into the planner to determine the next navigation direction . When the agent arrives at a navigation point, it assesses whether to halt its journey. If the decision is ‘No’, the process loops back to the memory update step; if ‘Yes’, the agent responds to the question.
3 Method
MemoryEQA can efficiently handle complex tasks that involve multiple targets across different regions. Our goal is to provide proper memory information as inputs to various modules, rather than using memory only at the time of answering.
3.1 Background
Typically, an EQA agent consists of four modules: a planner $\mathcal{F}_{P}$, a memory module $\mathcal{F}_{M}$, an answering module $\mathcal{F}_{A}$, and a stopping module $\mathcal{F}_{S}$. We define the observation sequence as $O=\{{o_{i}}\}_{i=0}^{N_{t}}$ and pose sequence as $P=\{p_{i},r_{i}\}_{i=0}^{N_{t}}$ received by the agent, where $p_{i}$ and $r_{i}$ is the $i$-th position and rotation, $N_{t}$ represents the total number of exploration steps. We use $\mathcal{M}=\{o_{i},p_{i},m_{i}\}_{i=0}^{N_{t}}$ represent the memory library. The usual EQA framework can be formalized as
[…]
where $q$ is the question, and $r$ is the response of VLMs. $\mathcal{F}_{S}(\cdot)$ is an indicator function: if the current observation can answer the question, it outputs 1; otherwise, it outputs 0. During the exploration process, the agent will determine whether it can answer the question based on the current observation. If $\mathcal{F}_{S}(o_{i})=0$, it will continue to explore to obtain the next observation. Otherwise, it will directly answer the question. However, this approach only utilizes historical information for the answering module $\mathcal{F}_{A}$, which may result in redundant or insufficient exploration, leading to suboptimal success rates.
3.2 Memory-Centric EQA Framework
We propose the memory-centric EQA (MemoryEQA) framework, as shown in Figure 1, which formulates the memory module $\mathcal{F}_{M}=\{f_{M}^{r},f_{M}^{u}\}$ consisting of a retrieval function ($f_{M}^{r}$) and an update function ($f_{M}^{u}$). This module converts the retrieved information into an appropriate form as input to each component. We define the memory information as $\mathcal{M}=\{m_{i}\}_{i=0}^{N_{s}}=\{f_{M}^{u}(o_{i})\}_{i=0}^{N_{s}}$, where $o_{i}$ is the image observation, and $m_{i}$ the natural language memory. MemoryEQA can be formulated as
[…]
where $q_{p}$, $q_{s}$, and $q_{a}$ denote the queries used to retrieve memory information required by the planner, stopping module, and answering module, respectively (details see the supplementary materials).
3.2.1 Planner
We extend the classical frontier-based exploration [yamauchi1997frontier] with memory guidance, which integrates top-$p$ frontier selection and VLM-based reasoning. The environment is represented as an occupancy grid map $\mathcal{G}=\{g_{i,j}\}$, where each cell is labeled as free, occupied, or unknown. A frontier cell is defined as a free cell adjacent to at least one unknown cell, and connected frontier cells are grouped into regions $\mathcal{D}=\{D_{1},D_{2},\cdots,D_{N}\}$. Each frontier $D_{i}$ is evaluated by a utility function $U(D_{i})=\lambda_{1}G(D_{i})-\lambda_{2}C(D_{i})$, where $G(\cdot)$ measures expected information gain and $C(\cdot)$ denotes navigation cost. $\lambda_{1}$ and $\lambda_{2}$ are weight coefficients used to balance the importance of information gain and navigation cost in target selection. Instead of selecting only the frontier with the highest utility, we retain the top-$p$ candidates to preserve diverse exploration directions:
$\mathcal{D}^{*}={\text{Top-}p}_{D_{i}\in\mathcal{D}}(U(D_{i})).$
Each selected frontier centroid is projected to pixel space using the camera model, from which a visual patch is extracted. For frontier candidates, a query $q_{p}$ is used to retrieve memory information providing relevant past observations and semantics. The VLM then evaluates the consistency between multiple directions of the current observation and retrieved memory information, producing relevance scores. We ultimately chose the exploration direction with the highest relevance score. This strategy enables the agent to balance efficient spatial coverage with semantic relevance, where frontier utility ensures broad exploration, and memory-augmented VLM reasoning prioritizes task-relevant goals.
3.2.2 Memory Module
The memory module has three mechanisms: storage, update, and retrieval. Storage and update are used to maintain the memory library, while retrieval is used to extract task-related memory information library for different modules. As shown in Figure 1, when other modules (planner, stopping module, and answering module) are called, the first step is to retrieve required memory information and then make further inferences based on the information.